# Advanced Vision Transformers and Open-Set Learning for Robust Mosquito Classification: A Novel Approach to Entomological Studies

### v1

@@ Abstract
Mosquito-related diseases pose a significant threat to global public health, necessitating efficient and accurate mosquito classification for effective surveillance and control. This work presents an innovative approach to mosquito classification by leveraging state-of-the-art vision transformers and open-set learning techniques. A novel framework has been introduced that integrates Transformer-based deep learning models with comprehensive data augmentation and preprocessing methods, enabling robust and precise identification of ten mosquito species. The Swin Transformer model achieves the best performance for traditional closed-set learning with 99.80% accuracy and 0.998 F1 score. The lightweight MobileViT technique attains an almost similar accuracy of 98.90% with significantly reduced parameters and model complexities. Next, the applied deep learning models’ adaptability and generalizability in a static environment have been enhanced by using new classes of data samples during the inference stage that have not been included in the training set. The proposed framework’s ability to handle unseen classes like insects similar to mosquitoes, even humans, through open-set learning further enhances its practical applicability employing the OpenMax technique and Weibull distribution. The traditional CNN model, Xception, outperforms the latest transformer with higher accuracy and F1 score for open-set learning. The study’s findings highlight the transformative potential of advanced deep-learning architectures in entomology, providing a strong groundwork for future research and development in mosquito surveillance and vector control. The implications of this work extend beyond mosquito classification, offering valuable insights for broader ecological and environmental monitoring applications.
@@ 2 Related work
The morphological characterization is important for accurate mosquito species classification. In recent times, significant work has been performed on AI-based automatic mosquito detection to aid taxonomic experts and public health professionals. For this reason, many researchers have shown their interest in classifying mosquitoes using deep learning approaches. Most of these used various pretrained classification and transformer models to classify mosquitoes through images and wingbeat sounds. Some recent articles on AI-based automatic mosquito detection have been briefly discussed in the following paragraphs.
Rustam et al. [6] used machine learning algorithms and innovative RIFS feature selection approaches to classify vector mosquito images for disease epidemiology. The authors classified two disease-spreading classes of mosquitoes, Aedes and Culex, using an open-source dataset from the IEEE data port. VGG16 model with the RIFS feature selection approach accomplished 99.2 accuracy with 0.986 F1 coefficient. A mosquito categorization scheme using transfer learning was developed by Isawasan et al. [7]. The authors classified two species of Aedes mosquito with a public database of 2400 images. They used five CNN models with transfer learning approaches (ResNet50, Xception, InceptionV3, MobileNetV2, and VGG16), and all the applied models achieved accuracies of more than 90 with VGG16 attaining a maximum accuracy of 98.25.
Siddiqui and Kayte [8] applied various deep learning frameworks for mosquito classification. They collected a dataset of six species of mosquitoes from website search engines employing web-scrapping techniques. The applied custom CNN and VGG16 models attained the maximum accuracies of 97.17 and 85.75, respectively. A comparative analysis of conventional and deep learning techniques for mosquito species categorization was investigated by Okayasu et al. [9]. The authors employed 14400 images from a public repository and three types of mosquito classes. The conventional SIFT method achieved the highest accuracy of 0.824. In contrast, the ResNet deep learning method had the highest of 0.955, demonstrating the effectiveness of deep classification in the categorization of mosquito species.
Lee et al. [10] investigated deep learning-integrated picture categorization for principal mosquito species in Korea. They used a dataset of 12 classes and eleven mosquito species comprising 7096 images. The Swin transformer hybrid model offered the best performance with an F1 score of 0.971 and 96.5 precision. Zhao et al. [11] initiated mosquito species identification utilizing various deep learning techniques. The dataset consists of seven genera and 17 species, with a total of 9900 mosquito images. The Swin transformer model accomplished the maximum accuracy, which is . YOLOv5 and DETR gave good competition to Swin transformers with an accuracy of and , respectively.
@@ 3.1.1 Formation of the Dataset
The comprehensive dataset of this work has been constructed by combining various open-source image datasets utilized in related mosquito classification articles, e.g., image dataset of Aedes and Culex mosquito species [17], mosquito-on-human-skin [18], dataset of vector mosquito images [19] and malaria vector mosquito images [20]. The proposed dataset contains images of ten species of mosquitoes: two species of Aedes, seven of Anopheles and one Culex type. The dataset contains a total of 219 images of Aedes aegypti, 500 images of Aedes albopictus, 64 images of Anopheles albimanus, 193 images of Anopheles arabiensis, 36 images of Anopheles atroparvus, 74 images of Anopheles coluzzi, 57 images of Anopheles farauti, 97 images of Anopheles freeborni, nine images of Anopheles stephensi, and 500 images of Culex quinquefasciatus. The images of the two classes, Aedes albopictus and Culex quinquefasciatus, contain both whole and smashed/crushed mosquitoes, which can be described as partial mosquitoes, resulting in only parts of the insect being visible or intact. Table 1 represents the exact numbers of images for various mosquito classes. A sample image of each class is shown in Figure 1. Aedes is represented as Ae, Anopheles as An, and Culex as Cx in the caption of the figure.
| Type | Total |
| Aedes aegypti | 219 |
| Aedes albopictus | 500 |
| Anopheles albimanus | 64 |
| Anopheles arabiensis | 193 |
| Anopheles atroparvus | 36 |
| Anopheles coluzzi | 74 |
| Anopheles farauti | 57 |
| Anopheles freeborni | 97 |
| Anopheles stephensi | 9 |
| Culex quinquefasciatus | 500 |
| Total | 1749 |
@@ 3.1.2 Dataset Preprocessing
The data preprocessing approach is implemented to enhance the number of images by adding slightly altered imprints of previously existing data or creating new artificial information from the current data. It functions as a regularizer and helps prevent overfitting when a model is being trained [21]. Oversampling in data analysis is strongly connected to data augmentation techniques. In this work, additional images have been created from the original picture using standard data augmentation techniques, including rotation (20 degrees), width and height shifts (0.20), shear (0.20), zoom (0.20), horizontal flips, and nearest fill mode using ImageDataGenerator framwrok [22]. Figure 2 displays the number of augmented images of each mosquito class.
@@ 3.2 Model Training
This subsection describes the applied vision transformer-based deep learning models with their model architectures, open-set learning implementation and corresponding performance metrics.
@@ 3.2.1 ResNet50
ResNet50 is a deep convolutional neural network with 50 layers, utilizing residual learning to reduce the vanishing gradient problem through shortcut connections [23]. Training and validation datasets were loaded using flow_from_directory and image_dataset_from_directory with image dimensions 224224 and batch size 32. Excluding the top layer, the ResNet50 model was initialized with pre-trained weights from ImageNet. A new top layer was added consisting of a flattened layer, a dense network with 512 units and ReLU activation, a dropout layer (0.50), and a final dense layer with softmax activation. The model was compiled using the Adam optimizer (learning rate 0.0001), sparse categorical cross-entropy loss and 20 epochs. It is interesting to mention that, the Ray Tune Python library has been utilized in this work to select the optimized hyperparameters discussed above.
@@ 3.2.2 Xception
Xception enhances the Inception deep learning architecture by substituting the standard Inception modules with depthwise separable convolutions, leading to improved efficiency and performance [24]. The Xception model was initialized with pre-trained weights from ImageNet, excluding the top layer. Various hyperparameters involving the top layer for Xception models are found to be almost similar to ResNet50.
@@ 3.2.3 Vision Transformer (vit-base-patch16-224)
The Vision Transformer (ViT) applies the Transformer architecture to image analysis by treating image patches as sequences. This method enables the model to capture long-range dependencies within the image effectively. The specific variant utilized in this study is vit_base_patch16_224, which processes images in patches of 1616 [25]. Each image was resized to 256256 pixels for image preprocessing, center-cropped to 224224 pixels, randomly horizontally flipped, and normalized. The ViT model was instantiated using the timm library with pre-trained weights and was further adapted to classify different mosquito species, aligning with the research objectives. The model was trained using the Adam optimizer with a learning rate of 0.0001 and cross-entropy loss.
@@ 3.2.4 Swin Transformer (Swin-B and Swin-S)
Swin Transformers achieve efficiency by focusing their self-attention mechanism on localized image regions that shift across processing layers [26]. This design allows the model to efficiently handle long-range dependencies with a manageable computational complexity [27]. The Swin Transformer models are transferred to the computational device to ensure efficient processing. These setups leverage the pre-trained models to accelerate training and enhance the model’s ability to generalize from the dataset used for classifying mosquito species. Both models were trained using the Adam optimizer with a learning rate of 0.0001 and cross-entropy loss. The Swin-S model was trained for five epochs, while the Swin-B model was trained for ten epochs, with a batch size of 32. The architecture of the Swin Transformer used in this work is illustrated in Figure 3.
@@ 3.2.5 Convolutional Vision Transformer (CVT-13)
The Convolutional Vision Transformer integrates convolutional layers with vision transformer components, resulting in a model that efficiently leverages local and global features for image classification [28]. The training images underwent additional augmentation techniques, such as random horizontal flips, to introduce variability and improve the model’s robustness to different orientations and lighting conditions. In contrast, the test images were only resized and center-cropped to maintain the integrity of the test conditions. The CVT-13 model was initialized with pre-trained weights from the Microsoft model hub and tailored for the classification of mosquito species. A feature extractor was configured with AutoFeatureExtractor.from_pretrained(’microsoft/cvt-13’, do_rescale=False), allowing direct use of the model’s capabilities without resizing the images to the size initially used during the training of the model. The model, CvtForImageClassification.from_pretrained(’microsoft/cvt-13’, num_labels=len(train_dataset.classes), ignore_mismatched_sizes=True), was then transferred to the computational device. This setup ensures compatibility with the dataset’s specific class sizes and feature requirements while allowing for efficient computation. The architecture of the CVT-13 model employed in this study is depicted in Figure 4.
@@ 3.2.6 MobileViT
MobileViT integrates the efficiency of MobileNets approaches with the representational power of Vision Transformers, resulting in a lightweight yet robust model suitable for mobile and resource-constrained environments [29]. Data was resized to 256256, center-cropped to 224224, randomly horizontally flipped, and normalized. The MobileViT model was created using the transformers library with pre-trained weights, modified for mosquito species classification. The model was trained using the Adam optimizer (learning rate 0.0001) and cross-entropy loss for ten epochs with batch size 32. The architecture of the MobileViT model applied in this work is shown in Figure 5.
@@ 3.3 Open-set Learning Implementation
The dataset for open-set learning has been prepared through the following steps:
OpenMax Dataset preparation: Construct a test set of 100 unknown images (unseen during the training step) – bee (33), butterfly (34), and fly (33).
Use a trained model to calculate the Mean Activation Vector (MAV) and fit the Weibull distribution.
Create a pickle file.
Use the pickle file to set an OpenMax threshold.
@@ 3.3.6 OpenMax Threshold:
3.4 Workflow
Figure 6 illustrates a workflow diagram for the proposed deep learning-based mosquito classification system involving data collection, preprocessing, model training, and evaluation. The process begins with data collection and merging various mosquito images, followed by data splitting and augmentation. It includes a decision point to check if the data is balanced, and if not, the data is balanced before compiling and training various models (ResNet50, Xception, ViT, Swin-S, Swin-B, CVT, and MobileViT). The training is followed by a check for the desired accuracy, with options to adjust the corresponding hyperparameters and re-evaluate the models. Once the desired accuracy is achieved, open-set learning is implemented employing the OpenMax dataset and Weibull distribution, and reports are generated.

### v2

@@ Abstract
Mosquito-related diseases pose a significant threat to global public health, necessitating efficient and accurate mosquito classification for effective surveillance and control. This work presents an innovative approach to mosquito classification by leveraging state-of-the-art vision transformers and open-set learning techniques. A novel framework has been introduced that integrates Transformer-based deep learning models with comprehensive data augmentation and preprocessing methods, enabling robust and precise identification of ten mosquito species. The Swin Transformer model achieves the best performance for traditional closed-set learning with 99.60% accuracy and 0.996 F1 score. The lightweight MobileViT technique attains an almost equivalent accuracy of 98.90% with significantly reduced parameters and model complexities. Next, the applied deep learning models’ adaptability and generalizability in a static environment have been enhanced by using new classes of data samples during the inference stage that have not been included in the training set. The proposed framework’s ability to handle unseen classes like insects similar to mosquitoes, even humans, through open-set learning further enhances its practical applicability employing the OpenMax technique and Weibull distribution. The traditional CNN model, Xception, outperforms the latest transformer with higher accuracy and F1 score for open-set learning. The study’s findings highlight the transformative potential of advanced deep-learning architectures in entomology, providing a strong groundwork for future research and development in mosquito surveillance and vector control. The implications of this work extend beyond mosquito classification, offering valuable insights for broader ecological and environmental monitoring applications.
Author Summary
Mosquitoes are major vectors for diseases like malaria, dengue, and Zika, causing widespread health challenges globally. Identifying mosquito species accurately is critical for controlling these diseases. Traditionally, mosquito classification is done manually, which is labor-intensive and time-consuming. To address this, we have developed an automated system that uses advanced deep learning models, specifically vision transformers, to classify mosquitoes into ten distinct species, including Aedes, Anopheles, and Culex, using a dataset of 5000 images. Our approach not only improves mosquito identification in static environments but also introduces a novel feature called open-set learning, enabling the system to detect and flag insects or species that were not part of the training dataset. This makes the system highly adaptable for real-world applications where unexpected or new mosquito species may appear. By combining these cutting-edge technologies, our work aims to support more effective mosquito surveillance, aiding public health initiatives and disease prevention efforts, particularly in regions most affected by mosquito-borne illnesses.
@@ 2 Related work
The morphological characterization is important for accurate mosquito species classification. In recent times, significant work has been performed on AI-based automatic mosquito detection to aid taxonomic experts and public health professionals. For this reason, many researchers have shown their interest in classifying mosquitoes using deep learning approaches. Most of these used various pretrained classification and transformer models to classify mosquitoes through images and wingbeat sounds. Some recent articles on AI-based automatic mosquito detection are briefly discussed in the following paragraphs.
Rustam et al. [6] used machine learning algorithms and innovative RIFS feature selection approaches to classify vector mosquito images for disease epidemiology. The authors classified two disease-spreading classes of mosquitoes, Aedes and Culex, using an open-source dataset from the IEEE data port. VGG16 model with the RIFS feature selection approach accomplished 99.2 accuracy with 0.986 F1 coefficient. A mosquito categorization scheme using transfer learning was developed by Isawasan et al. [7]. The authors classified two species of Aedes mosquito with a public database of 1205 images. They used five CNN models with transfer learning approaches (ResNet50, Xception, InceptionV3, MobileNetV2, and VGG16), and all the applied models achieved accuracies of more than 90 with VGG16 attaining a maximum accuracy of 98.25.
Siddiqui and Kayte [8] applied various deep learning frameworks for mosquito classification. They collected a dataset of six species of mosquitoes from website search engines employing web-scrapping techniques. After augmentation, the dataset contained a total of 5400 images. The applied custom CNN and VGG16 models attained the maximum accuracies of 97.17 and 85.75, respectively. A comparative analysis of conventional and deep learning techniques for mosquito species categorization was investigated by Okayasu et al. [9]. The authors employed 7200 images clicked through camera and three types of mosquito classes. The conventional SIFT method achieved the highest accuracy of 0.824. In contrast, the ResNet deep learning method had the highest of 0.955, demonstrating the effectiveness of deep classification in the categorization of mosquito species.
Lee et al. [10] investigated deep learning-integrated picture categorization for principal mosquito species in Korea. They used a dataset of 12 classes and eleven mosquito species comprising 7096 images after augmentation. The Swin transformer hybrid model offered the best performance with an F1 score of 0.971 and 96.5 precision. Zhao et al. [11] initiated mosquito species identification utilizing various deep learning techniques. The dataset consists of seven genera and 17 species, with a total of 9900 mosquito images after augmentation. The Swin transformer model accomplished the maximum accuracy, which is . YOLOv5 and DETR gave good competition to Swin transformers with an accuracy of and , respectively.
@@ 3.1.1 Formation of the Dataset
The comprehensive dataset of this work has been constructed by combining various open-source image datasets utilized in related mosquito classification articles, e.g., image dataset of Aedes and Culex mosquito species [17], mosquito-on-human-skin [18], dataset of vector mosquito images [19] and malaria vector mosquito images [20]. The proposed dataset contains images of ten species of mosquitoes: two species of Aedes, seven of Anopheles and one Culex type. The images of the two classes, Aedes albopictus and Culex quinquefasciatus, contain both whole and smashed/crushed mosquitoes, which can be described as partial mosquitoes, resulting in only parts of the insect being visible or intact. Table 1 represents the exact numbers of images for various mosquito classes. A sample image of each class is shown in Figure 1. Aedes is represented as Ae, Anopheles as An, and Culex as Cx in the caption of the figure.
Type | Total without Augmentation | Total with Augmentation |
Aedes aegypti | 219 | 500 |
Aedes albopictus | 500 | 500 |
Anopheles albimanus | 64 | 500 |
Anopheles arabiensis | 193 | 500 |
Anopheles atroparvus | 36 | 500 |
Anopheles coluzzi | 74 | 500 |
Anopheles farauti | 57 | 500 |
Anopheles freeborni | 97 | 500 |
Anopheles stephensi | 9 | 500 |
Culex quinquefasciatus | 500 | 500 |
Total | 1749 | 5000 |
@@ 3.1.2 Dataset Preprocessing
The data preprocessing approach was applied to expand the dataset by introducing slightly modified versions of existing images or generating new artificial data from the current images. This process acts as a regularizer, mitigating overfitting during model training [21]. Data augmentation techniques are strongly connected to oversampling in this context, allowing for an enhanced and balanced dataset. In this work, additional images were generated using standard augmentation techniques, including rotation (limited to 20 degrees), width and height shifts (0.20), shear (0.20), zoom (0.20), horizontal flips, and nearest fill mode via the ImageDataGenerator framework [22].
The choice to limit the rotation to 20 degrees was intentional to preserve the natural wing positions and antennae orientation of mosquitoes, as these anatomical features are essential for accurate species classification. Larger rotations or deformations could distort these key features, negatively affecting the model’s ability to distinguish between species. Additionally, the augmentation parameters such as shifts and zoom were set at 20% to introduce sufficient variation without significantly altering the mosquito’s structural integrity. This augmentation approach is widely considered a well-established standard in the field, balancing the need for diversity while maintaining the essential characteristics of the objects of interest. Figure 2 presents the number of augmented images for each mosquito class.
3.1.3 Dataset Splitting
After data preprocessing and augmentation, the comprehensive dataset of 5000 images was split into training, validation, and test sets. Initially, the dataset was divided into a training set and a test set using an 80:20 ratio, resulting in 4000 images in the training set and 1000 images in the test set. The 4000 training images were then further split into a training set and a validation set, using another 80:20 ratio. This resulted in 3200 images for the final training set and 800 images for the validation set.
Thus, the final dataset consisted of 3200 images for training, 800 images for validation, and 1000 images for testing. This stratified splitting ensured that each mosquito species was proportionally represented across all sets, maintaining class balance throughout. During the training phase, the training set (3200 images) was used to fit the models, while the validation set (800 images) was employed to monitor performance and fine-tune hyperparameters. After training was completed, the test set (1000 images) was utilized for the final evaluation of the model on unseen data.
-0.3in0in Dataset Split Number of Images Percentage (%) Description Training Set 3200 64% Used for model training Validation Set 800 16% Used for performance monitoring Test Set 1000 20% Used for final model evaluation Total 5000 100% Complete dataset
@@ 3.2 Model Training
This subsection describes the applied vision transformer-based deep learning models with their model architectures, open-set learning implementation and corresponding performance metrics. All experiments were conducted on Google Colab Pro, utilizing the NVIDIA L4 GPU for training, which offers a balanced performance for large models like Swin Transformers. The environment ran on Ubuntu Linux, providing a stable platform for deep learning tasks. We used Python along with several key libraries: TensorFlow and Keras for model training and evaluation, PyTorch and Hugging Face Transformers for Vision Transformer (ViT) models, and Timm for managing advanced transformer architectures like Swin-B and MobileViT. This combination allowed for efficient model implementation and optimization within the deep learning workflow.
@@ 3.2.1 ResNet50
ResNet50 is a deep convolutional neural network with 50 layers, leveraging residual learning to mitigate the vanishing gradient problem by using shortcut connections [23]. The model was fine-tuned using pre-trained ImageNet weights. A custom classification head was added, consisting of a dense layer with 512 units, ReLU activation, and a dropout layer to prevent overfitting. The final output layer employed softmax activation for classifying mosquito species. The model was trained using the Adam optimizer with a learning rate of 0.0001 and sparse categorical cross-entropy loss over 20 epochs. Hyperparameter optimization was conducted to improve the model’s performance manually before the model training began, based on best practices and fixed values.
@@ 3.2.2 Xception
Xception is an extension of the Inception architecture, replacing the standard Inception modules with depthwise separable convolutions, which enhances model efficiency and performance [24]. This model was fine-tuned using pre-trained weights from ImageNet. A new classification head was added similarly to ResNet50, and training followed the same optimization strategy.
@@ 3.2.3 Vision Transformer (vit-base-patch16-224)
The Vision Transformer (ViT) model utilizes a transformer architecture designed for image classification by treating image patches as sequences, capturing long-range dependencies in the image [25]. The specific variant used was vit_base_patch16_224, which splits images into 16×16 patches. The model was pre-trained on ImageNet and fine-tuned for mosquito classification. Image preprocessing involved resizing, center-cropping, and augmenting to improve robustness. The Adam optimizer with a learning rate of 0.0001 and cross-entropy loss was used for training.
@@ 3.2.4 Swin Transformer (Swin-B and Swin-S)
Swin Transformers efficiently handle image data by focusing attention on localized regions, shifting across layers [26, 27]. This design allows the model to capture both local and global image features. Swin-B and Swin-S models were fine-tuned using pre-trained ImageNet weights and adapted to the mosquito classification task. Both models used the Adam optimizer and cross-entropy loss. The Swin-S model was trained for five epochs, while Swin-B was trained for ten epochs, with a batch size of 32.
The architecture of the Swin Transformer used in this work is illustrated in Figure 3.
@@ 3.2.5 Convolutional Vision Transformer (CVT-13)
The Convolutional Vision Transformer (CVT) integrates convolutional layers with transformer components, enabling the model to effectively capture both local features and global context [28]. Data augmentation was applied to enhance variability and robustness during training. The model was pre-trained on ImageNet and fine-tuned for mosquito classification. The training images were resized, center-cropped, and augmented to maintain consistency, while the CVT-13 model was trained using similar optimization strategies as other models.
The architecture of the CVT-13 model employed in this study is depicted in Figure 4.
@@ 3.2.6 MobileViT
MobileViT combines the efficiency of MobileNets with the representational power of Vision Transformers, making it ideal for mobile and resource-constrained environments [29]. This lightweight model was pre-trained and fine-tuned for mosquito species classification. Image preprocessing included resizing, center-cropping, and applying random augmentations to improve generalization. The MobileViT model was trained using the Adam optimizer with a learning rate of 0.0001 and cross-entropy loss for ten epochs with a batch size of 32.
The architecture of the MobileViT model applied in this work is shown in Figure 5.
3.2.7 Evaluation Metrics
To comprehensively assess the performance of the trained models, several evaluation metrics were employed, including accuracy, precision, recall, F1-score, Receiver Operating Characteristic (ROC) curve, and confusion matrix. These metrics provide a detailed understanding of how well the models perform on different aspects of classification, such as overall correctness (accuracy), the ability to correctly identify positive instances (precision), the capability to find all relevant instances (recall), and the harmonic mean of precision and recall (F1-score). The ROC curve was used to analyze the trade-off between the true positive rate and the false positive rate at various threshold levels, while the confusion matrix offered insights into the distribution of true positives, true negatives, false positives, and false negatives. Importantly, all these evaluation metrics were calculated using the test set consisting of 1000 images, which was completely held out during the model training and validation phases. This ensures that the evaluation results reflect the models’ performance on unseen data, providing a robust measure of their generalization capabilities.
@@ 3.3 Open-set Learning Implementation
The dataset for open-set learning was prepared by constructing a new extended test set comprising a new class ”unknown” with the existing 10 classes. This new class has 100 unknown images that were not seen during the training phase. This test set includes images of bees (33 images), butterflies (34 images), and flies (33 images). To utilize the trained model, we calculated the Mean Activation Vector (MAV) for the known classes and fitted the Weibull distribution to model the tails of these activations. Finally, a threshold for the OpenMax layer was established based on the Weibull distribution, which helps in determining the likelihood of a sample belonging to an unknown class.
@@ 3.3.6 OpenMax Threshold:
3.4 Top-level overview of the proposed mosquito classification system.
Figure 6 illustrates the workflow for the proposed deep learning-based mosquito classification system. The process begins with the collection and merging of various mosquito image datasets. After the data is collected, it is split and augmented to increase variability and dataset size. At a decision point, the workflow checks if the data is balanced. If the data is unbalanced, balancing techniques are applied to ensure even class distribution. Once the data is balanced, various models, including ResNet50, Xception, Vision Transformer (ViT), Swin-S, Swin-B, Convolutional Vision Transformer (CVT), and MobileViT, are compiled and trained. The next step in the workflow is checking if the models have achieved the desired accuracy. If the performance does not meet the expectations, hyperparameters are adjusted, and the models are retrained. Once the desired accuracy is achieved, the workflow proceeds to implement open-set learning using the OpenMax layer and the Weibull distribution to handle unknown classes during inference. The final step in the process involves generating classification reports based on the trained model’s performance.